OpenStudy (anonymous):

Suppose that the integral of x-1 of f(t) dt = 5x^2 + 7x -3. Find f(x)

OpenStudy (anonymous):

\[\int\limits_{x}^{1} f(t) dt = 5x^2 + 7x - 3\] is that written correctly ?

OpenStudy (anonymous):

x is on the top

OpenStudy (anonymous):

if so, use the fact that \[\int\limits_{b}^{a} f(x) dx = -\int\limits_{a}^{b} f(x) dx\] and the fundamental theorem of calculus \[d/dx \int\limits_{a}^{x} f(t) dt = f(x)\]

OpenStudy (anonymous):

ah, in that case just use the fundamental theorem of calculus

OpenStudy (anonymous):

i.e., take the derivative of both sides with respect to x

OpenStudy (anonymous):

quick question, when writing the integral out, a should come before b? like how I put x-1, should it have been x-a?

OpenStudy (anonymous):

i read it as "from x to 1", so to be clear in writing it out you could say "from a to b"

OpenStudy (anonymous):

or more precisely, "from [lower limit] to [upper limit]"

OpenStudy (anonymous):

ok thanks! I was confused about that as well

OpenStudy (anonymous):

what you take the derivative with respect to x on both sides, what do you end up with ?